1. What number does not belong 1/16,1/4,1/6,1/2,1,2
2.A plane arrived Tuesday at 8:00 pm Wednesday it was 64 seconds late. How many seconds late would it be in a half hour?

Answer by stanbon(75887)

1. What number does not belong 1/16,1/4,1/6,1/2,1,2
All the denomators are powers of 2 except 6.
Ans: 1/6
----------------------------------
2.A plane arrived Tuesday at 8:00 pm Wednesday it was 64 seconds late. How many seconds late would it be in a half hour?
1/2 hour = (1/2)(3600 seconds) = 1800 seconds
----------------
Cheers,
Stan H.
--------------